If cos θ=7/25 , find the values of all T-ratios of θ.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given function: cos θ = 7/25

We know that cos θ = AB/AC = Base/Hypotenuse = 7/25

By Pythagoras Theorem:

AC^2 = AB^2 + BC^2
BC^2 = AC^2 - AB^2
= 625k^2–49k^2
= 576k^2

AB = 24k

Find other T-rations using their definitions:
Sin𝜃= BC/AC = 24/25

Tan𝜃= BC/AB = 24/7
cosec 𝜃= 1/sin𝜃= 25/24
sec𝜃= 1/cos𝜃= 25/7
cot𝜃= 1/tan𝜃= 7/24

Thanks